OpenStudy (mollykenney):

write each quadratic function in vertex for. Find the coordinates of the vertex and the equation of the axis of symmetry. f(x) = -1/2x^2

OpenStudy (bibby):

the axis of symmetry is \(\huge \frac{-b}{2a}\) where \(\large ax^2+bx+c\). Here b is 0 so we have an axis of symmetry of \(\huge \frac{-0}{2*\frac{-1}{2}}= \frac{0}{-1}\) or just 0 a look at the graph confirms this

OpenStudy (bibby):

similarly the vertex would be (0,0)

OpenStudy (mollykenney):

THANK YOU @bibby

OpenStudy (bibby):

np, yo